# Predicting the adhesion and delamination strength of carbon films on metals by high-throughput ab initio calculations

### v1

@@ Abstract
Diamond and diamond-like carbon (DLC) coatings are widely employed for their exceptional mechanical, thermal and chemical properties, but their industrial application is often limited by weak adhesion to metallic substrates. In this work, we employ a high-throughput ab initio approach to systematically investigate the adhesion of diamond/metal interfaces, combining a set of technologically relevant metals (Al, Ag, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W, Zn) with the C(111), C(111)-21 (Pandey reconstructed), C(110), C(100), that are most common in diamond and are representative of different types of bonds present in DLC. Thanks to our automated and accurate computational protocol for interface construction and characterization, databases are populated and relevant trends are identified on the effect of surface graphitization, ability to form carbides and metal reactivity on carbon film adhesion and delamination strength. Beyond capturing trends, our workflow yields predictive insights. Indeed, we found that adhesion energy scales with the geometric mean of the constituent surface energies, providing a simple descriptor for rapid screening; while comparing the work of separation with the metal’s cohesive energy anticipates the fracture location under tensile loading. A novel method based on analysis is introduced to identify when contact with a metal drives rehybridization of surface carbon from sp2 to sp3, the structural signature of improved resistance to delamination. These structural changes are mirrored by an electronic rearrangement at the interface, quantified by a charge-accumulation descriptor that strongly correlates with adhesion.
I INTRODUCTION
@@
In this work, we address this gap using an ab initio high-throughput approach, enabling a systematic evaluation and comparison of adhesion energies across technologically relevant diamond/metal interfaces. Furthermore, we propose a rigorous computational protocol for constructing the interface supercell. In contrast to the common practice of simply adapting the lattice of one component to the other, typically by straining the metallic substrate to match the diamond lattice often resulting in unphysical deformations, our approach employs the algorithm developed by Zur, which identifies the smallest common supercell between two crystalline surfaces while controlling lattice mismatch and angular distortion below user-defined limits [32]. A detailed description of the algorithm and its implementation is provided in Section II.
On the metal side, this study considers the most stable terminations of Al, Ag, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W and Zn; while four different low-index diamond surfaces are chosen to represent the different types of bonds in DLC coatings: the non-reconstructed C(111)(11) surface with dangling bond (DB) termination, its most stable reconstruction C(111)(21) called Pandey reconstruction, the C(110)(11) surface and the so-called Dimer reconstructed C(100)(21) termination. From now on, we will address them as C(111), C(111)Pd, C(110) and C(100), respectively. C(111) is the only one exhibiting a sp3 bond structure at the surface; while C(111) Pd, with its zig-zag surface chains and delocalized electrons, is representative of the sp2 aromatic or graphitic layer commonly observed on the diamond/DLC surface during rubbing. C(110) also forms zig-zag -bonded chains, but its stabilization and inertness is less pronounced than that of C(111)Pd; whereas C(100) stabilizes through surface dimers, resulting in a distorted sp2-like geometry [33, 34, 35].
The paper is organized as follows: in Section II, we describe the computational protocol implemented in our high-throughput software for the automated construction and characterization of diamond/metal interfaces. Section III presents and discusses the results, focusing first on the trends in adhesion energies across the various diamond terminations and metal substrates and on how these relate to the intrinsic properties of the constituents, such as their surface energy. Then, the fracture behavior of the interfaces is examined to assess their mechanical stability under tensile load. Finally, electronic and structural responses upon interface formation, such as charge accumulation and modification of diamond surface reconstruction, are addressed, providing a microscopic understanding of the adhesion mechanisms.
II METHODS
@@
A key feature of TribChem is its systematic storing of relevant physical results. In general, the database is a fundamental component in high-throughput workflows; however, in TribChem, beyond the internal FireWorks database used for tasks management, a new high-level one is created to store physical quantities returned by the workflow in an organized manner. In particular, we use MongoDB, a NoSQL database, which facilitates future data analysis and ensures that data are not recomputed.
@@
A schematic representation of the computational workflow is shown in Fig. 1. The first step consists in retrieving the input structures for the selected materials from the online Materials Project database [48] and storing them in our database. Both bulk structures are then optimized in parallel, by converging their lattice parameters via a fit with the Birch-Murnaghan equation of state [49] and optimizing the kinetic energy cutoff and k-point density mesh required by DFT. The higher values for energy cutoff and k-point density are chosen after all convergence loops have finished, so that the interfaces calculations will be both efficient and accurate. Subsequently, slabs are obtained by cutting the bulk along the selected directions and their thickness is converged with respect to the surface energy. The physical quantities obtained from the bulk and surface workflows for the selected metals and diamond terminations are reported in Tab.1; while optimized computational parameters are shown in Tab.S1. The diamond surfaces characterized by complex reconstructions, namely C(111)Pd and C(100), were constructed outside the automated workflow, but still using the bulk parameters optimized by TribChem.
@@
As already anticipated in Section I, the matching of the two slabs to construct the heterointerface is performed following the Zur algorithm [32], implemented in the MPInterfaces library, which constructs the smallest supercell possible while satisfying specific criteria on lattice dimensions and angles strains. Further details on the algorithm are thoroughly discussed in Ref.[46]. The magnitude of the applied strain is scaled according to the inverse of the bulk modulus of each material, meaning that materials with higher compressibility undergo greater deformation during interface construction. However, in this case, since diamond is significantly stiffer than the metals considered, the strain was applied exclusively to the metallic component. Since lattices with different symmetries must be accommodated within a single supercell, the strain on the metal is anisotropic in some cases. Fig.2 reports the maximum percentage strain, while average strain values are provided in Fig.S1. For all metals, the lattice strain is below 3, which significantly reduces the impact of mechanical stress on the interface electronic properties. Details on the transformations applied to the slab cells to construct the interface supercell, as automatically computed by TribChem, can be found in the Supporting Information (SI).
Finally, the tribological properties of the interface such as the Potential Energy Surface (PES), adhesion energy and charge transfer at the interface are calculated. The PES maps the interaction energy as a function of the relative lateral position of the two slabs. It is therefore obtained by relaxing the slabs in the direction normal to the interface plane for all the relative lateral shifts obtained by matching the two surfaces’ high symmetry points. Adhesion energy is then computed as:
namely by subtracting the total energies of the relaxed single slabs, and , from the total energy of the interface in the lateral position which corresponds to the PES minimum , divided by the supercell area . Electronic charge accumulation upon interface formation () is calculated as:
@@
III RESULTS AND DISCUSSION
Considering a subset of technologically relevant transition metals and including Al and Mg simple metals due to their importance in alloys, the most stable facets of Ag, Al, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W and Zn have been mated to four low-index diamond surfaces: C(111), C(111)Pd, C(110) and C(100). A total of 60 diamond/metal interfaces were optimized and their adhesion were calculated. This requires to sample the whole PES and identify the energy minimum, meaning that a total of 950 DFT calculations were performed. The atomic models of interfaces formed by the four diamond terminations in contact with representative metallic substrate are shown in Fig.4.
III.1 Adhesion energy
Adhesion energies, as calculated from Eq.1, are reported in Fig.3. In general, constituents with higher surface energies form an interface characterized by higher adhesion (bottom left corner of Fig.3), whereas diamond terminations adhere less to noble metals, such as Cu, Ag, Au, which possess completely filled d-shells.
When examining Fig.3 row by row, i.e. by varying the metal substrate while keeping the diamond termination fixed, it is possible to identify some outliers. In particular, Ti shows strong adhesion to all diamond terminations, despite its relatively low reactivity, quantified by its surface energy. This behaviour can be attributed to its pronounced carbide-forming ability compared to other metals (carbide formation enthalpy of TiC per Ti atom =–185.2 kJ/mol [18] and TiC formation energy of -0.808 eV/atom [48]). Conversely, when examining the values in Fig.3 column by column, i.e. at fixed metal substrate, diamond terminations with higher surface energy exhibit stronger adhesion across all metals. An interesting comparison can be made between C(111), which exposes surface sp3 carbons, and C(111)Pd, characterized by chains of sp2 hybridized carbon atoms. While previous work reported adhesion reduction at C/Cu due to surface graphitization [50], here we see that this behavior is general and that surface graphitization drastically reduces adhesion across all metals. This suggests that while growing the DLC film, for an effective coating that maximizes the adhesion to the substrate while reducing adhesion and friction with any countersurface, a gradient of the sp2/sp3 fraction should be built with larger sp2/sp3 ratio.
@@ III.1 Adhesion energy
III.2 Adhesion energies from intrinsic constituent properties
@@
The effectiveness of this relationship is further confirmed for diamond/metal interfaces, as illustrated in Fig.5, where is plotted against at fixed diamond termination to account for the different nature of surface bonds. and RMSE values along with the linear fit parameters are reported in Tab.2. In all cases, the relatively high values, lying in the range 70-79, indicate that the overall trend predicted by the model is robust; while deviations in slope and intercept from the ideal values of 1 and 0, respectively, highlight that predicting adhesion energy is a complex task, governed by the interplay of multiple factors, such as chemical and structural variations at the interface.
@@ III.3 Fracture location
Fig.6 shows the difference between the work of separation of the semiconductor/metal interface and interlayer cohesive energy of the metal. When C(111)Pd is involved, the fracture always occurs at the interface, reflecting the fact that surface graphitization of diamond leads to extremely weak bonds with the metal and, consequently, to very low adhesion, increasing the risk of delamination. For metals with low surface energy (/2), e.g., Mg, Zn, Al, fracture may occur within the metal itself, especially when in contact with the highly reactive C(111) surface. On the other hand, metals with partially filled d-shells (W, Cr, Mo, Fe) exhibit extremely high cohesion energies, which prevent failure from taking place within the metal phase.
III.4 Structural and electronic modifications of diamond reconstructions upon adhesion to metals
@@
The of the standalone C(111)Pd surface is shown in Fig.7(a). Peaks corresponding to the interatomic distances within the surface atoms are indicated by vertical dashed lines and illustrated in the adjacent schematic. Remaining peaks correspond to the bulk diamond structure. Here the focus is on the peak at 1.44 Å, which corresponds to the distances between atoms in the surface zig-zag chain. The area under a peak is directly related to the average number of atoms in the interval [, ] surrounding a given atom, i.e., the number of bonds per atom contributing to that peak, namely [58]:
where is the average atomic density (total number of atoms over cell volume). For C(111)Pd, the integral of the peak at 1.44 Å therefore provides a measure of the number of surface sp2 bonds per atom.
To assess the impact of metal contact on the surface graphitization, Fig.7(c) shows a qualitative comparison of the of the pristine C(111)Pd surface (plotted in the background with faded lines) with that of the same surface but within the interface geometries (black line). Different types of structural modifications are observed depending on the metal: in some cases, the peak intensity decreases without significant shift or distortion (Mo, W, Ti, Fe), while in others, dimerization (Cr, Al, Cu, Pt) or a shift in (Rh, V, Zn) is evident, reflecting a more pronounced perturbation of the surface reconstruction. Quantitatively, integrating the peak provides a measure of how many sp2 bonds remain after interaction with each metal. For example, the weak interaction with Au leaves the peak essentially unaltered, showing only minor broadening and indicating that all sp2 bonds are preserved. In contrast, Cr, which exhibits strong adhesion with C(111)Pd, induces partial dimerization of the chain and reduces the number of sp2 bonds by roughly 50. The adhesion energy of all C(111)Pd/metal interfaces as a function of the fraction of sp2 bonds preserved relative to the pristine surface is plotted in Fig.7(b). Less reactive metals, i.e., those with completely filled outer electron shells (Ag, Au, Zn, Mg), preserve the original number of sp2 bonds of the C(111)Pd surface, explaining their weak adhesion. On the contrary, more reactive metals are effectively able to perturb the surface sp2 network, leading to stronger adhesion. This behavior is highly relevant for coating applications, where metallic interlayers are commonly introduced to improve the bonding between DLC films and their substrates. In DLC systems with a high density of surface sp2 bonds, the interaction with the metal counteracts graphitization, responsible for the weak adhesion to the substrate, by promoting the rehybridization of surface carbon atoms.
@@
The rehybridization of the C(111)Pd surface chains is clearly reflected in the redistribution of electronic charge upon contact with the metal. As examples, we consider the interfaces with Cr and Ag, representing two extreme cases: the Cr interface exhibits dimerization of the surface chains and strong adhesion ( J/m2), while the Ag interface shows only a slight broadening of the peak, corresponding to weak adhesion ( J/m2). The volumetric charge difference, , shown in Fig.8(b) and (d), reveals that the charge rearrangement at the Cr interface is significantly more complex than at Ag. This is further quantified by the line profiles of along the direction, reported in Fig.8(a) and (c). In both cases the metal just accumulates charge in the interfacial region, while the surface carbon atoms reorganize their electronic charge density. Importantly, for the interface with Cr, this reorganization is far more pronounced: charge is removed at the surface carbon atom atomic plane where the C-C double bonds of the Pandey chain laid, moving towards the interface region, in line with the observed dimerization of the surface chains.
III.5 Charge density accumulation at the interface

### v2

# Insights into the adhesion and delamination strength of carbon films on metals by high-throughput ab initio calculations

@@ Abstract
Diamond and diamond-like carbon (DLC) coatings are widely employed for their exceptional mechanical, thermal and chemical properties, but their industrial application is often limited by weak adhesion to metallic substrates. In this work, we employ a high-throughput ab initio approach to systematically investigate the adhesion of diamond/metal interfaces, combining a set of technologically relevant metals (Al, Ag, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W, Zn) with the C(111), C(111)-21 (Pandey reconstructed), C(110), C(100) surfaces, that are most common in diamond and are representative of different types of bonds present in DLC. Thanks to our automated and accurate computational protocol for interface construction and characterization, databases are populated and relevant trends are identified on the effect of surface graphitization, ability to form carbides and metal reactivity on carbon film adhesion and delamination strength. Beyond capturing trends, our workflow yields predictive insights. Indeed, we found that adhesion energy scales with the geometric mean of the constituent surface energies, providing a simple descriptor for rapid screening; while comparing the work of separation with the metal’s cohesive energy anticipates the fracture location under tensile loading. A novel method based on the radial distribution function analysis is introduced to identify when contact with a metal drives rehybridization of surface carbon from sp2 to sp3, the structural signature of improved resistance to delamination. These structural changes are mirrored by an electronic rearrangement at the interface, quantified by a charge-redistribution descriptor that strongly correlates with adhesion.
1 INTRODUCTION
@@
In this work, we address this gap using an ab initio high-throughput approach, enabling a systematic evaluation and comparison of adhesion energies across technologically relevant diamond/metal interfaces. Furthermore, we propose a rigorous computational protocol for constructing the interface supercell. In contrast to the common practice of simply adapting the lattice of one component to the other, typically by straining the metallic substrate to match the diamond lattice often resulting in unphysical deformations, our approach employs the algorithm developed by Zur, which identifies the smallest common supercell between two crystalline surfaces while controlling lattice mismatch and angular distortion below user-defined limits [61]. A detailed description of the algorithm and its implementation is provided in Section 2.
On the metal side, this study considers the most stable low-index terminations of Al, Ag, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W and Zn, i.e. the crystallographic orientations with the lowest surface energies, since these facets are the most likely to be exposed under equilibrium conditions. We note that chemical modifications of metal surfaces are expected to alter the adhesion energy, because the interfacial bonding is strongly affected by the local surface chemistry, as shown in previous works [45, 10]. On the diamond side, four different low-index surfaces are chosen to represent the different types of bonds in DLC coatings: the non-reconstructed C(111)(11) surface with dangling bond (DB) termination, its most stable reconstruction C(111)(21) called Pandey reconstruction, the C(110)(11) surface and the so-called Dimer reconstructed C(100)(21) termination. From now on, we will address them as C(111), C(111)-rec, C(110) and C(100), respectively. C(111) is the only one exhibiting a sp3 bond structure at the surface; while C(111)-rec, with its zig-zag surface chains and delocalized electrons, is representative of the sp2 aromatic or graphitic layer commonly observed on the diamond/DLC surface during rubbing. C(110) also forms zig-zag -bonded chains, but its stabilization and inertness is less pronounced than that of C(111)-rec; whereas C(100) stabilizes through surface dimers, resulting in a distorted sp2-like geometry [11, 49, 3].
The paper is organized as follows: in Section 2, we describe the computational protocol implemented in our high-throughput software for the automated construction and characterization of diamond/metal interfaces. Section 3 presents and discusses the results, focusing first on the trends in adhesion energies across the various diamond terminations and metal substrates and on how these relate to the intrinsic properties of the constituents, such as their surface energy. Then, the fracture behavior of the interfaces is examined to assess their mechanical stability under tensile load. Finally, electronic and structural responses upon interface formation, such as charge redistribution and modification of diamond surface reconstruction, are addressed, providing a microscopic understanding of the adhesion mechanisms.
2 METHODS
@@
A key feature of TribChem is its systematic storing of relevant physical results. In general, the database is a fundamental component in high-throughput workflows; however, in TribChem, beyond the internal FireWorks database used for tasks management, a new high-level one is created to store physical quantities returned by the workflow in an organized manner. In particular, we use MongoDB, a NoSQL database.
@@
A schematic representation of the computational workflow is shown in Fig. 1. The first step consists in retrieving the input structures for the selected materials from the online Materials Project database [23] and storing them in our database. Both bulk structures are then optimized in parallel, by converging their lattice parameters via a fit with the Birch-Murnaghan equation of state [6] and optimizing the kinetic energy cutoff and k-point density mesh required by DFT. The higher values for energy cutoff and k-point density are chosen after all convergence loops have finished, so that the interfaces calculations will be both efficient and accurate. Subsequently, slabs are obtained by cutting the bulk along the selected directions and their thickness is converged with respect to the surface energy. The physical quantities obtained from the bulk and surface workflows for the selected metals and diamond terminations are reported in Tab.1; while optimized computational parameters are shown in Tab.S1. The diamond surfaces characterized by complex reconstructions, namely C(111)-rec and C(100), were constructed outside the automated workflow, but still using the bulk parameters optimized by TribChem.
@@
As already anticipated in Section 1, the matching of the two slabs to construct the heterointerface is performed following the Zur algorithm [61], implemented in the MPInterfaces library, which constructs the smallest supercell possible while satisfying specific criteria on lattice dimensions and angles strains. Further details on the algorithm are thoroughly discussed in Ref.[46]. The magnitude of the applied strain is scaled according to the inverse of the bulk modulus of each material, meaning that materials with higher compressibility undergo greater deformation during interface construction. However, in this case, since diamond is significantly stiffer than the metals considered, the strain was applied exclusively to the metallic component. Since lattices with different symmetries must be accommodated within a single supercell, the strain on the metal is anisotropic in some cases. Fig.2 reports the maximum percentage strain, while average strain values are provided in Fig.S1. For all metals, the lattice strain is below 3, which significantly reduces the impact of mechanical stress on the interface electronic properties. Details on the transformations applied to the slab cells to construct the interface supercell, as automatically computed by TribChem, can be found in the Supporting Information (SI).
Finally, the tribological properties of the interface such as the Potential Energy Surface (PES), adhesion energy and charge transfer at the interface are calculated. The PES maps the interaction energy as a function of the relative lateral position of the two slabs. It is therefore obtained by relaxing the slabs in the direction normal to the interface plane for all the relative lateral shifts obtained by matching the two surfaces’ high symmetry points. The resulting PES is then used to identify the minimum-energy registry. Starting from this configuration, a subsequent fully unconstrained relaxation (atomic displacements allowed in all directions) is performed. Its total energy, , is the one employed to compute the adhesion energy:
where and are the total energies of the isolated slabs, relaxed with the same settings, and is the interface supercell area. For all relaxation calculations, the electronic self-consistency threshold is eV and the ionic relaxation is converged when residual forces are below 0.015 eV/Å. Electronic charge redistribution upon interface formation () is calculated as:
@@
3 RESULTS AND DISCUSSION
Considering a subset of technologically relevant transition metals and including Al and Mg simple metals due to their importance in alloys, the most stable facets of Ag, Al, Au, Cr, Cu, Fe, Ir, Mg, Mo, Pt, Rh, Ti, V, W and Zn have been mated to four low-index diamond surfaces: C(111), C(111)-rec, C(110) and C(100). A total of 60 diamond/metal interfaces were optimized and their adhesion were calculated. This requires to sample the whole PES and identify the energy minimum, meaning that a total of 950 DFT calculations were performed. The atomic models of interfaces formed by the four diamond terminations in contact with representative metallic substrate are shown in Fig.3.
3.1 Adhesion energy
Adhesion energies, as calculated from Eq.1, are reported in Fig.4. In general, constituents with higher surface energies form an interface characterized by higher adhesion (lower left region of Fig.4). This comprises in particular C(110) and C(111) diamond terminations in contact with reactive transition metal surfaces, like the bcc(110) early transition metals W, Cr, Mo, V. By contrast, all diamond terminations adhere less to metals which possess completely filled d-shells, such as Cu, Ag, Au.
When examining the values in Fig.4 column by column, i.e. at fixed metal substrate, diamond terminations with higher surface energy exhibit stronger adhesion across all metals. An interesting comparison can be made between C(111), which exposes surface sp3 carbons, and C(111)-rec, characterized by chains of sp2 hybridized carbon atoms. While previous work reported adhesion reduction at C/Cu due to surface graphitization [10], here we see that this behavior is general and that surface graphitization drastically reduces adhesion across all metals. This suggests that while growing the DLC film, for an effective coating that maximizes the adhesion to the substrate while reducing adhesion and friction with any countersurface, a gradient of the sp2/sp3 fraction should be built with larger sp2/sp3 ratio. Conversely, when examining Fig.4 row by row, i.e. by varying the metal substrate while keeping the diamond termination fixed, it is possible to observe that metals with a partially occupied d-shell adhere most strongly, especially Ti, Cr and V. In particular, Ti shows strong adhesion to all diamond terminations, despite its relatively low reactivity, quantified by its surface energy. This behaviour can be attributed to its pronounced carbide-forming ability compared to other metals (carbide formation enthalpy of TiC per Ti atom =–185.2 kJ/mol [16] and TiC formation energy of -0.808 eV/atom [23]).
@@ 3.1 Adhesion energy
3.2 Adhesion energies from intrinsic constituent properties
@@
The effectiveness of this relationship is further confirmed for diamond/metal interfaces, as illustrated in Fig.5, where is plotted against at fixed diamond termination to account for the different nature of surface bonds. and RMSE values along with the linear fit parameters are reported in Tab.2. In all cases, the relatively high values, lying in the range 70-79, indicate that the overall trend predicted by the model is robust; while deviations in slope and intercept from the ideal values of 1 and 0, respectively, highlight that predicting adhesion energy is a complex task, governed by the interplay of multiple factors, such as chemical and structural variations at the interface. The correlation can be further improved by adopting a more advanced data driven strategy, such as the Sure Independence Screening and Sparsifying Operator (SISSO) algorithm [41]. This method has already been successfully applied to heterogeneous metallic interfaces to identify the best descriptors and the optimal mathematical expression fitting the training data, by combining multiple material properties like cohesive energies and bulk moduli [46]. Here, however, our goal is to provide a physical and easily interpretable message. For this reason, we focus on the term that carried the largest weight in the previous SISSO-derived expression, i.e. , and that also has the most immediate physical meaning and experimental accessibility, rather than presenting a more complex multi-parameter expression.
@@ 3.3 Fracture location
Fig.6 shows the difference between the work of separation of the semiconductor/metal interface and interlayer cohesive energy of the metal. When C(111)-rec is involved, the fracture always occurs at the interface, reflecting the fact that surface graphitization of diamond leads to extremely weak bonds with the metal and, consequently, to very low adhesion, increasing the risk of delamination. For metals with low surface energy (/2), e.g., Mg, Zn, Al, fracture may occur within the metal itself, especially when in contact with the highly reactive C(111) surface. On the other hand, metals with partially filled d-shells (W, Cr, Mo, Fe) exhibit extremely high cohesion energies, which prevent failure from taking place within the metal phase.
3.4 Structural and electronic modifications of diamond reconstructions upon adhesion to metals
@@
The of the standalone C(111)-rec surface is shown in Fig.7(a). Peaks corresponding to the interatomic distances within the surface atoms are indicated by vertical dashed lines and illustrated in the adjacent schematic. Remaining peaks correspond to the bulk diamond structure. Here the focus is on the peak at 1.44 Å, which corresponds to the distances between atoms in the surface zig-zag chain. The area under a peak is directly related to the average number of atoms in the interval [, ] surrounding a given atom, i.e., the number of bonds per atom contributing to that peak, namely [24]:
where is the average atomic density (total number of atoms over cell volume). For C(111)-rec, the integral of the peak at 1.44 Å therefore provides a measure of the number of surface sp2 bonds per atom.
To assess the impact of metal contact on the surface graphitization, Fig.7(b) shows a qualitative comparison of the of the pristine C(111)-rec surface (plotted in the background with faded lines) with that of the same surface but within the interface geometries (black line). Different types of structural modifications are observed depending on the metal: in some cases, the peak intensity decreases without significant shift or distortion (Mo, W, Ti, Fe, Ir), while in others, dimerization (Cr, Al, Cu, Pt, Mg) or a shift in (Rh, V, Zn) is evident, reflecting a more pronounced perturbation of the surface reconstruction. Quantitatively, integrating the peak provides a measure of how many sp2 bonds remain after interaction with each metal. For example, the weak interaction with Au leaves the peak essentially unaltered, showing only minor broadening and indicating that all sp2 bonds are preserved. In contrast, Cr, which exhibits strong adhesion with C(111)-rec, induces partial dimerization of the chain and reduces the number of sp2 bonds by roughly 50. The adhesion energy of all C(111)-rec/metal interfaces as a function of the fraction of sp2 bonds preserved relative to the pristine surface is plotted in Fig.7(c). Two characteristic regimes emerge, highlighted by the insets: weakly interacting metals, i.e., those with completely filled outer electron shells (Ag, Au, Zn), preserve essentially the original number of sp2 bonds of the C(111)-rec surface, consistently with their weak adhesion; conversely, more reactive metals are effectively able to perturb the surface sp2 network, leading to stronger adhesion. In particular, carbon-reactive/carbide-forming metals (Ti, Cr, Mo, W, V, Fe) preserve only about half of the original sp2 bonds. This behavior is highly relevant for coating applications, where metallic interlayers are commonly introduced to improve the bonding between DLC films and their substrates. In DLC systems with a high density of surface sp2 bonds, the interaction with the metal counteracts graphitization, responsible for the weak adhesion to the substrate, by promoting the rehybridization of surface carbon atoms.
@@
Although no change in reconstruction is expected for the unreconstructed C(111) termination upon contact (its primary response being the saturation of its highly reactive dangling bonds through C–metal bond formation), we extended the analysis to C(111) and to the metallic counterparts as a general tool to track slab structural evolution upon interface formation (see SI). In C(111) the surface bilayer sp3 C–C peak is mostly only reduced/broadened, with the strongest perturbations for Cr, Ti, V; for the metals, the largest peak changes are observed for the more strongly adhering transition metals (e.g., V, Ti, Mo, Cr).
@@
The rehybridization of the C(111)-rec surface chains is clearly reflected in the redistribution of electronic charge upon contact with the metal. As examples, we consider the interfaces with Cr and Ag, representing two extreme cases: the Cr interface exhibits dimerization of the surface chains and strong adhesion ( J/m2), while the Ag interface shows only a slight broadening of the peak, corresponding to weak adhesion ( J/m2). The volumetric charge difference, , shown in Fig.8(a) and (b), reveals that the charge rearrangement at the Cr interface is significantly more complex than at Ag. This is further quantified by the line profiles of along the direction, reported in Fig.8(c) and (d). In both cases the metal just accumulates charge in the interfacial region, while the surface carbon atoms reorganize their electronic charge density. Importantly, for the interface with Cr, this reorganization is far more pronounced: charge is removed at the surface carbon atom atomic plane where the C-C double bonds of the Pandey chain laid, moving towards the interface region, in line with the observed dimerization of the surface chains.
3.5 Charge density redistribution at the interface